The product of 2 more than a number and 10 is 36 more than 8 times the number. What is the number?
Answer by checkley77(12844)

(X+2)*10=8X+36
10X+20=8X+36
10X-8X=38-20
2X=18
X=9 ANS.
PROOF:
(9+2)*10=8*9+36
11*19=72+36
209=209